# Ladder Symmetry: The Necessary and Sufficient Condition for Vanishing Love Numbers

### v1

@@ I Introduction
The concept was introduced in the Newtonian setting by Love [Love:1909], who studied the deformation of elastic bodies under tidal forces. In the relativistic context, TLNs were later extended to neutron stars [Hinderer:2007mb, Damour:2009vw, Binnington:2009bb, Landry:2015cva]. This extension has profound astrophysical significance because TLNs provide a direct avenue to probe the equation of state of ultra-dense nuclear matter. Gravitational wave observations are particularly sensitive to tidal effects during the late inspiral phase of binary neutron star mergers, where the accumulation of finite-size corrections leaves observable signatures on the waveform. The landmark detection of GW170817 demonstrated this potential by delivering constraints on neutron star radii and on the nuclear equation of state [Abbott:2017xzu, Abbott:2018exr, Dietrich:2017aum].
@@
In contrast, black holes in four-dimensional and asymptotically flat general relativity (GR) display the remarkable property that their static TLNs vanish identically. This universality was first established for Schwarzschild black holes [Damour:2009vw, Binnington:2009bb] and later extended to the rotating Kerr family [Chia:2020yla, LeTiec:2020spy]. The absence of tidal deformability has become a defining feature of classical black holes, motivating its use as a null test of GR. Any observation of a non-vanishing TLN in connection with a black hole would constitute direct evidence for new physics beyond Einstein’s theory [Cardoso:2016rao, Pani:2015nua, Poisson:2014book].
@@
This observation naturally leads to the question of whether the vanishing of TLNs is simply an accidental property of black holes or whether it arises from a deeper underlying principle. A strong candidate for such a principle is the so-called Ladder symmetry, a hidden algebraic structure that appears in the perturbation equations of certain black hole spacetimes [Hui:2021vcv, Sharma:2024hlz]. Owing to this symmetry, the perturbation Hamiltonian for scalar, vector, and gravitational fluctuations in both Reissner–Nordström and Kerr backgrounds can be decomposed in terms of raising and lowering operators in close analogy with the harmonic oscillator. Within this framework, the ground state solution carries zero tidal response, and repeated application of the raising operator ensures the vanishing of TLNs for all higher modes.
@@
Building on this idea, ref. [Sharma:2024hlz] carried out a systematic study of Ladder symmetries in static and stationary black hole spacetimes beyond Kerr. A particularly useful framework is provided by the parametrization introduced by Konoplya, Rezzolla, and Zhidenko (the KRZ class), which offers a systematic and theory-agnostic description of deviations from Kerr while retaining desirable separability properties [Konoplya:2016jvv, Konoplya:2018arm]. By analyzing scalar perturbations in these geometries, [Sharma:2024hlz] identified a subclass of KRZ spacetimes that admit Ladder symmetry and consequently established that scalar TLNs vanish for the entire subclass. This finding demonstrated a clear correspondence between Ladder symmetry and the absence of tidal deformability in a wide family of black hole geometries.
@@
The present work aims to settle this question. We show that for static spherically symmetric black holes as well as rotating black holes within the KRZ class, Ladder symmetry is both a necessary and a sufficient condition for the vanishing of static TLNs. To demonstrate this, we employ the parameterized TLN framework developed in [Cardoso:2018ptl, Katagiri:2023umb] in combination with the most general Ladder-symmetric spacetimes classified in [Sharma:2024hlz]. By perturbing away from Ladder-symmetric backgrounds and analyzing the induced modifications to the effective perturbation potential, we establish that any departure from Ladder symmetry generically produces non-vanishing TLNs. This completes the logical equivalence and firmly identifies Ladder symmetry as the fundamental principle behind the no Love property of classical black holes.
@@ II Setup
Subsequently, in Sec. II.2, we revisit the results of Ref. [Sharma:2024hlz]. These results establish the most general forms of Ladder-symmetric spacetimes belonging to the Konoplya–Rezzolla–Zhidenko (KRZ) parametrization, both in the context of static, spherically symmetric black hole geometries and for stationary, rotating, axisymmetric black hole solutions. For the present analysis, we shall also demonstrate that it is sufficient to formulate the problem directly at the level of the perturbation (wave) equation, rather than starting from the full metric description. This observation significantly simplifies our calculations while retaining complete generality.
Finally, in Sec. II.3, we present the strategy underlying our approach. Using the Ladder-symmetric metrics described in Sec. II.2 as the background spacetimes, we will systematically quantify deviations from them within the parametrized framework introduced in Sec. II.1. This will provide a consistent and unified setting for our subsequent analysis of TLNs across different classes of black hole geometries.
@@ II.1 The Parametrized Formalism of calculating TLNs: A Brief Review
In [Cardoso:2019mqo, McManus:2019ulj, Katagiri:2023umb], the authors develop a parametrized formalism to compute various quantities of interest on spacetimes perturbatively deviated from known background solutions in a theory-agnostic manner. In [Cardoso:2019mqo, McManus:2019ulj], this formalism was employed for QNMs, while in [Katagiri:2023umb], it was used to derive tidal response in various scenarios. In [Kobayashi:2025swn], the authors used the same formalism to derive the tidal dissipation numbers for non-rotating black holes.
Here, we review the results of [Katagiri:2023umb] and discuss in Sec. II.3 how we adapt them to our analysis. The key assumptions underlying their approach are the following: (1) the background is spherically symmetric; (2) deviations from the background metric are small; and (3) no coupling exists between different perturbation sectors. Within this setup, the standard Zerilli/Regge–Wheeler and spin-$s$ perturbation equations are deformed by introducing small, linear power-law corrections to the effective potentials. These corrections, parameterized by coefficients, $\alpha_{j}^{\pm/s}$, capture the possible deviations from the background geometry.
Consider a static scalar ($\omega=0,~s=0$) field $\Xi\,(r,\theta,\varphi)=\chi_{\ell}(r)Y_{\ell m}(\theta,\varphi)$, where $Y_{\ell m}(\theta,\varphi)$ are the spherical harmonics of order $(\ell,m)$. The master equation governing these scalar perturbations on the Schwarzschild background is given by the radial equation of motion of $\chi_{\ell}(r)$ transformed into the Fourier domain,
@@
In such cases, the prefactor of the linear response term $(r_{+}/r)^{\ell}$ includes a logarithmic function dependent on $r$, and the constant numbers $K_{\ell}$’s are regarded as the fundamental parameters of a given system, which are independent of the scale measured, and then evaluated as “Running TLNs”. For Schwarzschild black holes, it is well-known that the horizon regularity condition enforces its static TLNs to vanish for all kinds of perturbations.
@@
Here, $j=0,1,2$ corrections are excluded from the analyses as these would be unboundedly large, and it is assumed that $|\alpha_{j}|<<r_{+}^{2}|V_{\ell}(r\to r_{+})|$ that suffices to ensure the smallness of all the other corrections. The linearity of the problem allows us to extract the physical tidal responses order by order in the small parameters $\alpha_{j}$’s by defining a “basis” set for TLNs. We can solve eq. (5) perturbatively by expanding $\chi_{\ell}$ up to the first order in $\alpha_{j}$ as
@@
with $e_{\ell,\,j}$’s as the “basis” set for TLNs, as termed in [Katagiri:2023umb]. Similarly, we can introduce a “basis” set $d_{\ell,\,j}$’s for the coefficient of running TLNs as
@@ II.2 Background Spacetimes with Ladder Symmetry
In [Sharma:2024hlz], the authors employ the presence of Ladder symmetries to impose constraints on the black hole spacetime metric. Here, we revisit the Ladder-symmetric solutions obtained therein, focusing on their role as background spacetimes within the parametrized framework. Furthermore, we will show that for the analyses and results presented in the following sections, it is sufficient to adopt the wave equation as the starting point, rather than the explicit form of the metric.
The most general form of the metric for a static, spherically symmetric four-dimensional spacetime where Ladder symmetry exists can be written as [Sharma:2024hlz]
@@
In asymptotically flat four-dimensional spacetimes, the most general class of stationary axisymmetric rotating black hole solutions where the scalar wave equation splits into radial and angular parts in the Boyer-Lindquist-type coordinates is denoted by the Konoplya-Rezzolla-Zhidenko (KRZ) class of metrics [Konoplya:2016jvv, Konoplya:2018arm]. The KRZ class includes well-known rotating black hole solutions like the Kerr, Kerr-Newman, and Kerr-Sen solutions.
The most general Ladder-symmetric form of the KRZ metric can be expressed as [Sharma:2024hlz]
@@
Now, depending on the choice of the residual gauge-fixing in the KRZ class of metrics, one can end up with different ways of quantifying the deviation from $f_{0}(r)$ in the metric. As shown in Sec. III, one such possible choice leads to having two different independent functions $f_{t}(r)=-g_{tt}$ and $f_{r}(r)=g_{rr}^{-1}$ in the wave equation. This way of encoding the deviations from a background metric can also be found in the EFT-corrected metrics studied in [Cardoso:2018ptl] and [Barbosa:2025uau]. Without loss of generality, a different choice of the gauge studied in [Sharma:2024hlz] can be implemented that leads to having only one independent function $\Delta(r)$ in the wave equation. In the following two sections III and IV, we will work with the choice used in [Sharma:2024hlz], such that all deviations from the background will be expressed as subleading terms in $\Delta(r)$. We will treat the metric and the resultant wave equation as our starting point and demonstrate how the parametrized formalism can be implemented in both these cases with two horizons in the background geometry.
@@ II.3 Strategy and Algorithm
We start with the metric of a generic stationary, rotating black hole belonging to the KRZ class [Konoplya:2018arm]
@@
Note that, if one were to show that there can exist metrics where TLNs vanish without the existence of Ladder-symmetry, then one would have to show that, besides the log-running contributions, the non-running contributions in the TLNs must vanish as well. Therefore, for our purpose, we could utilize these non-running contributions as well. But we will not do so for the following reasons. Firstly, in cases where the parametrized corrections to the effective potential are finite in number, there exists a maximum value of $\ell$ beyond which there will be no non-running contributions in the TLNs. Thus, even if we consider these contributions, in most cases, we will only have a finite number of such constraints on top of the infinite number of constraints obtained from the running contributions alone. Furthermore, as discussed in [Katagiri:2023umb], when the log-running contributions in the TLNs vanish identically, there can be ambiguities in reading off the pure tidal response terms from the resultant expressions, which are not resolved in the parametrized formalism. Using only the running TLNs allows us to bypass this subtlety of non-running TLNs. Therefore, we will employ only the running contributions to the TLNs to derive our main results.
@@ III Parametrized Love numbers for Static Spherically Symmetric Metrics
We have still not fixed the residual gauge freedom in eq. (22), and this would be a good point to shed light on some of the possible choices. The analysis of [Cardoso:2018ptl] corresponds to choosing $R_{\Sigma}=1$. For the general rotating KRZ solution, this choice corresponds to the Boyer-Lindquist coordinates for the Kerr metric. An alternative choice of this residual gauge is to fix $R_{B}=1$, which corresponds to Schwarzschild-like coordinates in the $a\to 0$ limit. From eq. 18, it is clear that setting either $a=0$ or $m=0$ in $V_{KRZ}$ yields equations that are schematically identical (although, the $m=0$ case still has an implicit dependence on $a$ through $R_{+}$ and $R_{-}$ in $f_{KRZ}$). This similarity implies that $m=0$ perturbations on a rotating KRZ background can be treated on the same footing as those in the $a\to 0$ (non-rotating) limit. Hence, in the subsequent calculations presented here, we use the $R_{B}=1$ gauge choice, which is consistent for both the static spherically symmetric and the stationary rotating KRZ metrics.
@@
In the subsequent calculations, we have omitted the $\ell$ from the subscripts of $\chi$ or $\Phi$ for brevity of notation. Following the algorithm in [Katagiri:2023umb] to cast the wave equation into a parametrized form, one can redefine the field $\chi(r)$ as
@@ V Conclusions and Discussions
The result extends naturally beyond the KRZ class. In more general or non-separable spacetimes, one may expect generalized Ladder-like structures that approximately suppress, but do not eliminate, tidal responses. Recent works like [Berens:2025jfs, Berens:2025okm] study the vanishing of tidal response for higher-spin perturbations in different dimensions using similar symmetries. It would be interesting to investigate whether, in those setups as well, the existence of these symmetries is a necessary criterion for the vanishing of TLNs. The vanishing of tidal response at a full nonlinear order has also been reported in [Parra-Martinez:2025bcu] for spherically symmetric black holes in asymptotically flat general relativity using worldline EFT. It would be interesting to use such worldline EFT calculations in our framework to understand the existence of running static scalar TLNs on spacetimes perturbatively deviated from Ladder-symmetric KRZ. In [Chakraborty:2025zyb], the authors have shown that the TLNs of a black hole perturbed by fermionic fields are zero. It would be interesting to probe it from a symmetry perspective and check whether it is possible to show that there never exists a “Ladder-like” symmetry in such cases, which would then correspond to non-vanishing TLNs for all perturbation modes. It would also be of interest to move beyond the KRZ parametrization to explore the manifestation of Love symmetry in more general stationary black hole spacetimes. Such an endeavor may also demand a deeper generalization of the very structure of the Love symmetry itself.

### v2

@@ I Introduction
The concept was introduced in the Newtonian setting by Love [33], who studied the deformation of elastic bodies under tidal forces. In the relativistic context, TLNs were later extended to neutron stars [23, 18, 10, 31]. This extension has profound astrophysical significance because TLNs provide a direct avenue to probe the equation of state of ultra-dense nuclear matter. Gravitational wave observations are particularly sensitive to tidal effects during the late inspiral phase of binary neutron star mergers, where the accumulation of finite-size corrections leaves observable signatures on the waveform. The landmark detection of GW170817 demonstrated this potential by delivering constraints on neutron star radii and on the nuclear equation of state [1, 2, 19].
@@
In contrast, black holes in four-dimensional and asymptotically flat general relativity (GR) display the remarkable property that their static TLNs vanish identically. This universality was first established for Schwarzschild black holes [18, 10] and later extended to the rotating Kerr family [17, 32]. Recent studies have further clarified this picture by systematically analyzing both static and dynamical tidal Love numbers of Schwarzschild and Kerr black holes in gravitational, scalar, and electromagnetic tidal fields, showing that while static Love numbers vanish, dynamical tidal responses are generically nonzero for rotating black holes. [6, 8, 7]. [^1] By contrast, for lower-dimensional black holes such as the BTZ solution, both static and dynamical scalar Love numbers are generically nonvanishing [9]. The absence of tidal deformability has become a defining feature of classical black holes, motivating its use as a null test of GR with asymptotically flat boundary conditions.  [^2] The role of asymptotic boundary conditions in the vanishing of tidal Love numbers is highlighted in [35, 21]. Also, [26, 16] addresses the instability of tidal Love numbers against environmental perturbations. Any observation of a non-vanishing TLN in connection with a black hole would constitute direct evidence for new physics beyond vacuum Einstein’s theory [12, 37, 39].
@@
This observation naturally leads to the question of whether the vanishing of static TLNs is simply an accidental property of black holes or whether it arises from a deeper underlying principle. A strong candidate for such a principle is the so-called Ladder symmetry, a hidden algebraic structure that appears in the perturbation equations of certain black hole spacetimes [24, 41]. Owing to this symmetry, the perturbation Hamiltonian for scalar, vector, and gravitational fluctuations in both Reissner–Nordström and Kerr backgrounds can be decomposed in terms of raising and lowering operators in close analogy with the harmonic oscillator. Within this framework, the ground state solution carries zero tidal response, and repeated application of the raising operator ensures the vanishing of TLNs for all higher modes.
@@
Building on this idea, ref. [41] carried out a systematic study of Ladder symmetries in static and stationary black hole spacetimes beyond Kerr. A particularly useful framework is provided by the parametrization introduced by Konoplya, Rezzolla, and Zhidenko (the KRZ class), which offers a systematic and theory-agnostic description of deviations from Kerr while retaining desirable separability properties [28]. The KRZ parametrization was introduced as a model-independent framework for describing black hole spacetimes in the strong-field regime, motivated by the increasing precision of gravitational-wave and electromagnetic observations [30]. As present and forthcoming experiments probe regions ever closer to the event horizon, it becomes necessary to adopt a systematic description of possible deviations from the Kerr geometry without committing to a specific theory of gravity. The KRZ framework addresses this need by providing a unified parametrization that captures strong-gravity effects while preserving key physical properties of black hole spacetimes, such as asymptotic flatness, the existence of a regular horizon, and, in the rotating case, the separability of perturbation equations. Beyond tidal Love numbers, the KRZ parametrization has been applied to a variety of strong-field phenomena, including quasinormal modes [29, 11], geodesic motion, and accretion-disk observables [22, 36, 43], as well as black hole shadows [42] and lensing properties [20], and has therefore been widely used in phenomenological tests of the Kerr-like geometry. By analyzing scalar perturbations in these geometries, [41] identified a subclass of KRZ spacetimes that admit Ladder symmetry and consequently established that scalar TLNs vanish for the entire subclass. This finding demonstrated a clear correspondence between Ladder symmetry and the absence of tidal deformability in a wide family of black hole geometries.
@@
The present work aims to settle this question. We show that for static spherically symmetric black holes as well as rotating black holes within the KRZ class, Ladder symmetry is both a necessary and a sufficient condition for the vanishing of static TLNs. To demonstrate this, we employ the parametrized TLN framework developed in [14, 25] in combination with the most general Ladder-symmetric spacetimes classified in [41]. By perturbing away from Ladder-symmetric backgrounds and analyzing the induced modifications to the effective perturbation potential, we establish that any departure from Ladder symmetry generically produces non-vanishing TLNs. This completes the logical equivalence and firmly identifies Ladder symmetry as the fundamental principle behind the no Love property of classical black holes.
@@ II Setup
Subsequently, in Sec. II.2, we revisit the results of Ref. [41]. These results establish the most general forms of Ladder-symmetric spacetimes belonging to the Konoplya–Rezzolla–Zhidenko (KRZ) parametrization, both in the context of static, spherically symmetric black hole geometries and for stationary, rotating, axisymmetric black hole solutions. For the present analysis, we shall also demonstrate that it is sufficient to formulate the problem directly at the level of the perturbation (wave) equation, rather than starting from the full metric description. This observation significantly simplifies our calculations while retaining complete generality.
Finally, in Sec. II.3, we present the strategy underlying our approach. Using the Ladder-symmetric metrics described in Sec. II.2 as the background spacetimes, we will systematically quantify deviations from them within the parametrized framework introduced in Sec. II.1.  [^3] In the static case, both the ladder-symmetric backgrounds considered here and the metrics obtained by perturbing away from ladder symmetry fall within the general class of static, spherically symmetric spacetimes described by the Rezzolla–Zhidenko parametrization [40]. Moreover, when the KRZ framework is restricted to the non-spinning limit and separability of the Klein–Gordon equation is imposed, it reduces to the static Rezzolla–Zhidenko form up to trivial gauge choices. This will provide a consistent and unified setting for our subsequent analysis of TLNs across different classes of black hole geometries.
@@ II.1 The Parametrized Formalism of calculating TLNs: A Brief Review
In [13, 34, 25], the authors develop a parametrized formalism to compute various quantities of interest on spacetimes perturbatively deviated from known background solutions in a theory-agnostic manner. In [13, 34], this formalism was employed for QNMs, while in [25], it was used to derive tidal response in various scenarios. In [27], the authors used the same formalism to derive the tidal dissipation numbers for non-rotating black holes.
Here, we review the results of [25] and discuss in Sec. II.3 how we adapt them to our analysis. The key assumptions underlying their approach are the following: (1) the background is spherically symmetric; (2) deviations from the background metric are small; and (3) no coupling exists between different perturbation sectors. Within this setup, the standard Zerilli/Regge–Wheeler and spin-$s$ perturbation equations are deformed by introducing small, linear power-law corrections to the effective potentials. These corrections, parametrized by coefficients, $\alpha_{j}^{\pm/s}$, capture the possible deviations from the background geometry.
Consider a static scalar ($\omega=0,~s=0$) field $\Xi\,(r,\theta,\varphi)=\frac{\chi_{\ell}(r)}{r}Y_{\ell m}(\theta,\varphi)$, where $Y_{\ell m}(\theta,\varphi)$ are the spherical harmonics of order $(\ell,m)$. The master equation governing these scalar perturbations on the Schwarzschild background is given by the radial equation of motion of $\chi_{\ell}(r)$ transformed into the Fourier domain,
@@
In such cases, the prefactor of the linear response term $(r_{+}/r)^{\ell}$ includes a logarithmic function dependent on $r$, and the constant numbers $K_{\ell}$’s are regarded as the fundamental parameters of a given system, which are independent of the scale measured, and then evaluated as “Running TLNs". For Schwarzschild black holes, it is well-known that the horizon regularity condition enforces its static TLNs to vanish for all kinds of perturbations.
@@
Here, $j=0,1,2$ corrections are excluded from the analyses as these would be unboundedly large, and it is assumed that $|\alpha_{j}|<<r_{+}^{2}|V_{\ell}(r\to r_{+})|$ that suffices to ensure the smallness of all the other corrections. The linearity of the problem allows us to extract the physical tidal responses order by order in the small parameters $\alpha_{j}$’s by defining a “basis" set for TLNs. We can solve eq. (5) perturbatively by expanding $\chi_{\ell}$ up to the first order in $\alpha_{j}$ as
@@
with $e_{\ell,\,j}$’s as the “basis” set for TLNs, as termed in [25]. Similarly, we can introduce a “basis” set $d_{\ell,\,j}$’s for the coefficient of running TLNs as
@@ II.2 Background Spacetimes with Ladder Symmetry
In [41], the authors employ the presence of Ladder symmetries to impose constraints on the black hole spacetime metric. Here, we revisit the Ladder-symmetric solutions obtained therein, focusing on their role as background spacetimes within the parametrized framework. Furthermore, we will show that for the analyses and results presented in the following sections, it is sufficient to adopt the wave equation as the starting point, rather than the explicit form of the metric.
The most general form of the metric for a static, spherically symmetric four-dimensional spacetime where Ladder symmetry exists can be written as [41]
@@
In asymptotically flat four-dimensional spacetimes, the most general class of stationary axisymmetric rotating black hole solutions where the scalar wave equation splits into radial and angular parts in the Boyer-Lindquist-type coordinates is denoted by the Konoplya-Rezzolla-Zhidenko (KRZ) class of metrics [30, 28]. The KRZ class includes well-known rotating black hole solutions like the Kerr, Kerr-Newman, and Kerr-Sen solutions.
The most general Ladder-symmetric form of the KRZ metric can be expressed as [41]
@@
Now, depending on the choice of the residual gauge-fixing in the KRZ class of metrics, one can end up with different ways of quantifying the deviation from $f_{0}(r)$ in the metric. As shown in Sec. III, one such possible choice leads to having two different independent functions $f_{t}(r)=-g_{tt}$ and $f_{r}(r)=g_{rr}^{-1}$ in the wave equation. This way of encoding the deviations from a background metric can also be found in the EFT-corrected metrics studied in [14] and [3]. Without loss of generality, a different choice of the gauge studied in [41] can be implemented that leads to having only one independent function $\Delta(r)$ in the wave equation. In the following two sections III and IV, we will work with the choice used in [41], such that all deviations from the background will be expressed as subleading terms in $\Delta(r)$. We will treat the metric and the resultant wave equation as our starting point and demonstrate how the parametrized formalism can be implemented in both these cases with two horizons in the background geometry.
@@ II.3 Strategy and Algorithm
We start with the metric of a generic stationary, rotating black hole belonging to the KRZ class [28]
@@
Note that, if one were to show that there can exist metrics where TLNs vanish without the existence of Ladder-symmetry, then one would have to show that, besides the log-running contributions, the non-running contributions in the TLNs must vanish as well. Therefore, for our purpose, we could utilize these non-running contributions as well. But we will not do so for the following reasons. Firstly, in cases where the parametrized corrections to the effective potential are finite in number, there exists a maximum value of $\ell$ beyond which there will be no non-running contributions in the TLNs. Thus, even if we consider these contributions, in most cases, we will only have a finite number of such constraints on top of the infinite number of constraints obtained from the running contributions alone. Furthermore, as discussed in [25], when the log-running contributions in the TLNs vanish identically, there can be ambiguities in reading off the pure tidal response terms from the resultant expressions, which are not resolved in the parametrized formalism. Using only the running TLNs allows us to bypass this subtlety of non-running TLNs. Therefore, we will employ only the running contributions to the TLNs to derive our main results.
@@ III Parametrized Love numbers for Static Spherically Symmetric Metrics
We have still not fixed the residual gauge freedom in eq. (22), and this would be a good point to shed light on some of the possible choices. The analysis of [14] corresponds to choosing $R_{\Sigma}=1$. For the general rotating KRZ solution, this choice corresponds to the Boyer-Lindquist coordinates for the Kerr metric. An alternative choice of this residual gauge is to fix $R_{B}=1$, which corresponds to Schwarzschild-like coordinates in the $a\to 0$ limit. From eq. 18, it is clear that setting either $a=0$ or $m=0$ in $V_{KRZ}$ yields equations that are schematically identical (although, the $m=0$ case still has an implicit dependence on $a$ through $R_{+}$ and $R_{-}$ in $f_{KRZ}$). This similarity implies that $m=0$ perturbations on a rotating KRZ background can be treated on the same footing as those in the $a\to 0$ (non-rotating) limit. Hence, in the subsequent calculations presented here, we use the $R_{B}=1$ gauge choice, which is consistent for both the static spherically symmetric and the stationary rotating KRZ metrics.
@@
In the subsequent calculations, we have omitted the $\ell$ from the subscripts of $\chi$ or $\Phi$ for brevity of notation. Following the algorithm in [25] to cast the wave equation into a parametrized form, one can redefine the field $\chi(r)$ as
@@ V Conclusions and Discussions
The result extends naturally beyond the KRZ class. In more general or non-separable spacetimes, one may expect generalized Ladder-like structures that approximately suppress, but do not eliminate, tidal responses. Recent works like [5, 4] study the vanishing of tidal response for higher-spin perturbations in different dimensions using similar symmetries. It would be interesting to investigate whether, in those setups as well, the existence of these symmetries is a necessary criterion for the vanishing of TLNs. The vanishing of tidal response at a full nonlinear order has also been reported in [38] for spherically symmetric black holes in asymptotically flat general relativity using worldline EFT. It would be interesting to use such worldline EFT calculations in our framework to understand the existence of running static scalar TLNs on spacetimes perturbatively deviated from Ladder-symmetric KRZ. In [15], the authors have shown that the TLNs of a black hole perturbed by fermionic fields are zero. It would be interesting to probe it from a symmetry perspective and check whether it is possible to show that there never exists a “Ladder-like" symmetry in such cases, which would then correspond to non-vanishing TLNs for all perturbation modes. It would also be of interest to move beyond the KRZ parametrization to explore the manifestation of Love symmetry in more general stationary black hole spacetimes. Such an endeavor may also demand a deeper generalization of the very structure of the Love symmetry itself.